OpenStudy (anonymous):

Rectangle A has an area of 16 − x2. Rectangle B has an area of x2 + 2x − 24. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

jimthompson5910 (jim_thompson5910):

Ratio of A to B (16-x^2)/(x^2+2x-24) From here, factor the numerator and denominator

jimthompson5910 (jim_thompson5910):

and cancel anything you can

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

What does 16-x^2 factor to?

OpenStudy (anonymous):

um i dont know honestly :'(

jimthompson5910 (jim_thompson5910):

16 - x^2 -x^2 + 16 -(x^2 - 16) -(x^2 - 4^2) -(x - 4)(x+4)

jimthompson5910 (jim_thompson5910):

I'm taking advantage of the difference of squares rule

jimthompson5910 (jim_thompson5910):

What do you get when you factor x^2+2x-24

OpenStudy (anonymous):

um okkay?

OpenStudy (anonymous):

? yah im stuck...i cant follow the same pattern as you just did cuz it longer... :/ i think

jimthompson5910 (jim_thompson5910):

x^2+2x-24 = (x-4)(x+6) since -4+6 = 2 and -4*6 = -24

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhh okkayyy

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{16 - x^2}{x^2+2x-24}\] becomes \[\Large \frac{-(x - 4)(x+4)}{(x-4)(x+6)}\]

OpenStudy (anonymous):

is that the simplest ratio? yes right?

jimthompson5910 (jim_thompson5910):

no, you need to cancel and simplify

OpenStudy (anonymous):

cancel what??

jimthompson5910 (jim_thompson5910):

notice how there's a "x-4" in common

OpenStudy (anonymous):

yes i do

jimthompson5910 (jim_thompson5910):

they cancel

jimthompson5910 (jim_thompson5910):

\[\Large \frac{-(x - 4)(x+4)}{(x-4)(x+6)}\] \[\Large \frac{-\cancel{(x - 4)}(x+4)}{\cancel{(x - 4)}(x+6)}\]

OpenStudy (anonymous):

but i tought they could only cancel if they were opposite? does the - sign include in that?

jimthompson5910 (jim_thompson5910):

no, the - out front is a -1 out front

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

which is a separate factor

jimthompson5910 (jim_thompson5910):

so we're left with \[\Large \frac{-(x+4)}{x+6}\] \[\Large \frac{-x-4}{x+6}\]

OpenStudy (anonymous):

so thats my anwser...

jimthompson5910 (jim_thompson5910):

yes, that's your ratio

OpenStudy (anonymous):

thank you so much! im like struggling in this module..:/

jimthompson5910 (jim_thompson5910):

just keep at it and you'll get it eventually

OpenStudy (anonymous):

@crististar